1. A copper wire is carrying 5.0 A of current. Given that the electric field between the ends of the cable has a value of 0.75 V/m. Answer the following questions: (50 points) . Find the current density in the wire. .Find the drift speed of the mobile charges. For this you will need the density of copper, D 8960 0.063 kg/mol. Assume each atom of iron contributes kg/m3 and the molar mass of copper, m with one electron. . If a potential difference of 20.0 Volts is applied, find the resistance on the cable. . Using the potential difference applied, find the length of the cable. Additionally, find the value of the cross sectional area of the cable.
Answer #1

1. J = E / rho


J = (0.75)/(1.72 x 10^-8)

J = 4.36 x 10^7 A/m^2


---------------------------------


mole in 1m^3 = (8960 )/(0.063) = 1.422 x 10^5 mol /m^3


free electron density = (1.422 x 10^5)(6.022 x 10^23)

= 8.565 x 10^28 electrons/m^3

Vd = J / n e

= (4.36 x 10^7) / (1.6 x 10^-19 x 8.565 x 10^28)

= 3.18 x 10^-3 m/s


----------------------------------------


R = V / I = 20 / 5 = 4 ohm


---------------------------

V = E L

20 = 0.75 L

L = 26.7 m

and J = I / A

4.36 x 10^7 = 5 / A

A = 1.15 x 10^-7 m^2